# The Nondecreasing Rank

### v1

@@ 8 ND Rank One and Two
If ${\mathbf{T}}\in\mathcal{C}(\times_{j=1}^{k}\mathcal{P}_{j})$ is a monotone, rank-one tensor, then $\mathbf{T}$} has an ND rank of one.
@@
As ${\mathbf{T}}\in\mathcal{N}_{<\infty}$ we have that ${\mathbf{T}}=\sum_{i=1}^{r}{\mathbf{a}}^{(i)}\otimes{\mathbf{b}}^{(i)}$, ${\mathbf{a}}^{(i)}\in\mathcal{C}(\mathcal{P}_{1})$, ${\mathbf{b}}^{(i)}\in\mathcal{C}(\mathcal{P}_{2})$, and ${\mathbf{T}}{\mathbf{h}}=\sum_{i=1}^{r}({\mathbf{h}}^{\intercal}{\mathbf{b}}^{(i)}){\mathbf{a}}^{(i)}\in\mathcal{C}(\mathcal{P}_{1})$ for all ${\mathbf{h}}\in\mathcal{C}(\mathcal{P}_{2})^{*}$. Similarly, ${\mathbf{h}}^{\intercal}{\mathbf{T}}\in\mathcal{C}(\mathcal{P}_{2})$ for all ${\mathbf{h}}\in\mathcal{C}(\mathcal{P}_{1})^{*}$. As $\mathcal{C}(\mathcal{P}_{2})^{*}$ is full-dimensional ${\mathbf{T}}(\mathcal{C}(\mathcal{P}_{2})^{*})\subseteq\mathcal{C}(\mathcal{P}_{1})$ is a two-dimensional cone. Let ${\mathbf{v}}^{(1)},{\mathbf{v}}^{(2)}$ be the extremal rays of the two-dimensional cone $\text{col}({\mathbf{T}})\cap\mathcal{C}(\mathcal{P}_{1})$. There exist ${\mathbf{w}}^{(1)},{\mathbf{w}}^{(2)}$ such that ${\mathbf{T}}={\mathbf{v}}^{(1)}\otimes{\mathbf{w}}^{(1)}+{\mathbf{v}}^{(2)}\otimes{\mathbf{w}}^{(2)}$ and it remains to show that ${\mathbf{w}}^{(i)}\in\mathcal{C}(\mathcal{P}_{2})$ for $i=1,2$. There must exist a face of $\mathcal{C}(\mathcal{P}_{1})$ that contains ${\mathbf{v}}^{(1)}$ but not ${\mathbf{v}}^{(2)}$, since otherwise $\lambda{\mathbf{v}}^{(1)}+(1-\lambda){\mathbf{v}}^{(2)}$ would be contained in $\mathcal{C}(\mathcal{P}_{1})$ for a small enough $\lambda>1$, contradicting the assumption that ${\mathbf{v}}^{(1)}$ was extremal in $\text{col}({\mathbf{T}})\cap\mathcal{C}(\mathcal{P}_{1})$. Let ${\mathbf{h}}\in\mathcal{C}(\mathcal{P}_{1})^{*}$ be a dual vector corresponding to the hyperplane that supports this face. Then ${\mathbf{h}}^{\intercal}{\mathbf{T}}=({\mathbf{h}}^{\intercal}{\mathbf{v}}^{(1)}){\mathbf{w}}^{(1)}+({\mathbf{h}}^{\intercal}{\mathbf{v}}^{(2)}){\mathbf{w}}^{(2)}=({\mathbf{h}}^{\intercal}{\mathbf{v}}^{(2)}){\mathbf{w}}^{(2)}\in\mathcal{C}(\mathcal{P}_{2})$, implying ${\mathbf{w}}^{(2)}\in\mathcal{C}(\mathcal{P}_{2})$ because ${\mathbf{h}}^{\intercal}{\mathbf{v}}^{(2)}>0$. The same argument shows that ${\mathbf{w}}^{(1)}\in\mathcal{C}(\mathcal{P}_{2})$. ∎
@@
A ramification of the above theorems is that defining equations for the semialgebraic sets [11, Def 2.1.4] of ND rank-one tensors and ND rank-two matrices are known. In the former case, a tensor $\mathbf{T}$} is in $\mathcal{N}_{\leq 1}$ if and only if it is monotonic and if it satisfies the determinantal equations for a rank-one tensor [33, Sec 3.4.1]. In the latter case, a matrix $\mathbf{T}$} is in $\mathcal{N}_{\leq 2}$ if and only if every $3\times 3$ minor of $\mathbf{T}$} vanishes and if $\mathbf{T}$} satisfies the inequalities for the $\mathcal{H}$-representation of $\mathcal{N}_{<\infty}$ described in Sections 5 and 6. Semialgebraic conditions under which a tensor of rank two also has a nonnegative rank of two are put forward in [3]. Whenever all order cones $\mathcal{C}(\mathcal{P}_{j})$ are simplicial such conditions can be easily translated, via Theorem 4, into conditions for a tensor to have an ND rank of two.
@@ 9 ND Border Rank Equals ND Rank
Due to the presence of noise, in many applications it is unlikely that a data tensor has an exact, low-rank, ND factorization. Instead a solution to an optimization problem of the form $\text{argmin}_{\bm{\theta}\in\mathcal{N}_{\leq r}}D({\mathbf{T}},\bm{\theta})$ is sought, where $D$ is a divergence that measures the discrepancy between the observed data tensor $\mathbf{T}$} and an approximating, low ND rank tensor $\bm{\theta}$. To ensure that a solution to this optimization problem exists it is important that the set $\mathcal{N}_{\leq r}$ be closed. The set of tensors with real tensor rank at most $r$ is not closed, which necessitates the introduction of the concept of border rank [33, Sec 2.4.5]. It is shown in this section, extending the proof for nonnegative tensor ranks provided in [37], that $\mathcal{N}_{\leq r}$ is closed and there is no need for the additional notion of border rank.
@@
Assume that $\bm{\theta}^{(n)}=\sum_{i=1}^{r}\lambda_{i}^{(n)}\otimes_{j=1}^{k}{\mathbf{v}}_{ij}^{(n)}$ is a sequence of ND rank $r$ tensors with $\|{\mathbf{T}}-\bm{\theta}^{(n)}\|_{F}\rightarrow\inf_{\bm{\theta}\in\mathcal{N}_{\leq r}}{\inf}\|{\mathbf{T}}-\bm{\theta}\|_{F}$. The ${\mathbf{v}}_{ij}^{(n)}\in\mathcal{C}(\mathcal{P}_{j})$ are assumed to be scaled so that $\|{\mathbf{v}}_{ij}^{(n)}\|_{2}=1$ and $\lambda_{i}^{(n)}\geq 0$. It is claimed that there exists an $M$ such that $\sup_{n}\lambda_{i}^{(n)}\leq M$ for all $i$. To show this, note that $\|{\mathbf{T}}-\bm{\theta}^{(n)}\|_{F}\geq\|\bm{\theta}^{(n)}\|_{F}-\|{\mathbf{T}}\|_{F}$ and
@@
where the last inequality follows because $\langle{\mathbf{v}}_{ij}^{(n)},{\mathbf{v}}_{i^{\prime}j}^{(n)}\rangle\geq 0$ for every $i,i^{\prime},j$ as $\mathcal{C}(\mathcal{P}_{j})\subseteq\mathbb{R}_{+}^{p_{j}}$. If $\lambda_{i}^{(n)}\rightarrow\infty$ for any $i$ then $\|{\mathbf{T}}-\bm{\theta}^{(n)}\|_{F}\rightarrow\infty$, a contradiction. It follows that there exists a convergent subsequence with $\lambda_{i}^{(n_{m})}\rightarrow\lambda_{i}$ and ${\mathbf{v}}_{ij}^{(n_{m})}\rightarrow{\mathbf{v}}_{ij}\in\mathcal{C}(\mathcal{P}_{j})$, with the latter inclusion following from $\mathcal{C}(\mathcal{P}_{j})$ being closed. Thus, $\bm{\theta}^{(n_{m})}\rightarrow\bm{\theta}=\sum_{i=1}^{r}\lambda_{i}\otimes_{j=1}^{k}{\mathbf{v}}_{ij}\in\mathcal{N}_{\leq r}$, proving that a solution to the specified optimization problems exists. To obtain the first statement, if $\mathbf{T}$} is in the closure of $\mathcal{N}_{\leq r}$, then $\inf_{\bm{\theta}\in\mathcal{N}_{\leq r}}\|{\mathbf{T}}-\bm{\theta}\|_{F}=0$. As this infimum can only be attained at $\bm{\theta}={\mathbf{T}}$, it follows that ${\mathbf{T}}=\bm{\theta}\in\mathcal{N}_{\leq r}$. ∎
@@ 10 Finding Low ND Rank Approximations
Solving the optimization problem $\inf_{\bm{\theta}\in\mathcal{N}_{\leq r}}\|{\mathbf{T}}-\bm{\theta}\|_{F}$ introduced in the previous section is a simple method for finding a low ND rank approximation to a data tensor $\mathbf{T}$}. However, when the entries of $\mathbf{T}$} consist of count data or positive data the Frobenius norm objective function may not be the most appropriate criteria to minimize. We examine a likelihood-based approach for estimating the mean of $\mathbf{T}$} in this section. It is assumed that the entries $T_{i_{1}\ldots i_{k}}$ are independently sampled from an exponential family of distributions with mean parameter [13, Ch 3] $E(T_{i_{1}\ldots i_{k}})=\theta_{i_{1}\ldots i_{k}}$ for every $(i_{1},\ldots,i_{k})\in\times_{j=1}^{k}[p_{j}]$, where $E(\cdot)$ is the expected value. The tensor $\bm{\theta}$ is constrained to have an ND rank of at most $r$ and is estimated by maximizing the observed likelihood
@@
Different distributional assumptions about $T_{i_{1}\ldots i_{k}}$ lead to different choices of $p(T_{i_{1}\ldots i_{k}}|\theta_{i_{1}\ldots i_{k}})$ and hence different optimization problems. Of primary interest are the following standard distributions:
@@
    If $T_{i_{1}\ldots i_{k}}\in\mathbb{R}$ then we may assume that $T_{i_{1}\ldots i_{k}}\sim\mathcal{N}(\theta_{i_{1}\ldots i_{k}},1)$ are independent Gaussians. The corresponding optimization problem is
@@
These four optimization problems represent special cases of $\beta$-divergences, that are commonly applied to NMF problems [24], where $\beta=2,1,0$ in $1.$ $3.$ and $4.$. Similar to Proposition 7.2 in [37], which shows that optimal nonnegative approximations exist for Bregman divergences, the next result shows that this also holds for nondecreasing factorizations, but does not require any additional constraints on $\bm{\theta}$.
@@
This follows for 1. by Theorem 13. For the remaining three problems it is seen that there exists an $M$ where any minimizer must be contained in the region $\mathcal{S}=\{\bm{\theta}:\theta_{i_{1}\ldots i_{k}}\leq M,\forall i_{1},\ldots,i_{k}\}$, as the objective functions in 3. and 4. diverge to $\infty$ whenever $\theta_{i_{1}\ldots i_{k}}\rightarrow\infty$. As $\mathcal{S}\cap\mathcal{N}_{\leq r}$ is compact by Theorem 13 and the objective functions are continuous on this region, existence of the solutions follows. Here the objective functions are extended continuously to take values in $\mathbb{R}\cup\{\infty\}$ with $T_{i_{1}\ldots i_{k}}\log(0)\coloneq\infty$ and $T_{i_{1}\ldots i_{k}}/0\coloneq\infty$ when $T_{i_{1}\ldots i_{k}}\neq 0$. ∎
@@
The Perron-Frobenius theorem [7, Thm 3.2] that states that nonnegative, square matrices have nonnegative eigenvectors, has been extended to nonnegative, cubical tensors in [36, Thm 1]. The next result is reminiscent of the Perron-Frobenius theorem, but is instead concerned with the singular vectors and values of a matrix rather than the eigenvectors and eigenvalues.
@@
If every fibre ${\mathbf{T}}_{i_{1}\ldots i_{j-1}\bullet i_{j+1}\ldots i_{k}}$ is in $\mathcal{C}(\mathcal{P}_{j})$ then the best rank-one approximation to $\mathbf{T}$} with respect to any of the log-likelihoods mentioned in this section is equal to the best ND rank-one approximation:
$\underset{\bm{\theta}:\mathrm{rank}(\bm{\theta})=1}{\mathrm{argmax}}\;p({\mathbf{T}}|\bm{\theta})=\underset{\bm{\theta}\in\mathcal{N}_{\leq 1}}{\mathrm{argmax}}\;\;p({\mathbf{T}}|\bm{\theta}).$
@@
Under the Gaussian likelihood, upon taking a gradient with respect to ${\mathbf{v}}^{(j)}$, the first-order conditions, without any monotonicity constraint, for $\bm{\theta}=\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}$ to be optimal are
@@
We may assume that any optimal rank-one approximation has ${\mathbf{v}}^{(j)}\geq{\mathbf{0}}$ as the tensor $\max(\otimes_{j=1}^{k}{\mathbf{v}}^{(j)},{\mathbf{0}})=\otimes_{j=1}^{k}\max({\mathbf{v}}^{(j)},{\mathbf{0}})$, where the maximum is taken elementwise, is closer to $\mathbf{T}$} than $\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}$. It is also clear that the optimal ${\mathbf{v}}^{(j)}$ is non-zero when $\mathbf{T}$} is non-zero. Equation (10) shows that ${\mathbf{v}}^{(j)}$ must be a conic combination of the monotone fibres ${\mathbf{T}}_{i_{1}\ldots i_{j-1}\bullet i_{j+1}\ldots i_{k}}$, and so it must itself be in $\mathcal{C}(\mathcal{P}_{j})$. When $k=2$ and ${\mathbf{v}}^{(1)},{\mathbf{v}}^{(2)}$ are any choice of the first left and right singular vectors by the Eckhart-Young theorem ${\mathbf{v}}^{(1)}\otimes{\mathbf{v}}^{(2)}$ is a best rank-one approximation of $\mathbf{T}$}. Therefore, by the above derivation, ${\mathbf{v}}^{(j)}\in\mathcal{C}(\mathcal{P}_{j})$, up to a possible sign change. If the largest singular value was repeated then we have pairs ${\mathbf{v}}^{(j)},{\mathbf{w}}^{(j)}$ of left ($j=1$) and right ($j=2$) singular vectors. As $\mathcal{C}(\mathcal{P}_{1})$ is proper, the subspace $\text{span}({\mathbf{v}}^{(1)},{\mathbf{w}}^{(1)})$ contains a unit norm vector ${\mathbf{z}}^{(1)}\notin\pm\mathcal{C}(\mathcal{P}_{1})$. Defining the $p_{j}\times 2$ matrix ${\mathbf{R}}^{(j)}=[{\mathbf{v}}^{(j)}|{\mathbf{w}}^{(j)}]$ consider the rotation matrix $\mathbf{V}$} constructed so that the first column of ${\mathbf{R}}^{(1)}{\mathbf{V}}$ is equal to ${\mathbf{z}}^{(1)}$. Set ${\mathbf{z}}^{(2)}$ to be the first column of ${\mathbf{R}}^{(2)}{\mathbf{V}}$. We have that ${\mathbf{R}}^{(1)}({\mathbf{R}}^{(2)})^{\intercal}={\mathbf{R}}^{(1)}{\mathbf{V}}{\mathbf{V}}^{\intercal}({\mathbf{R}}^{(2)})^{\intercal}$ is a best rank-two approximation of $\mathbf{T}$} and thus ${\mathbf{z}}^{(1)}\otimes{\mathbf{z}}^{(2)}$ is a best rank-one approximation of $\mathbf{T}$}. This contradicts the above fact that ${\mathbf{z}}^{(1)}\in\pm\mathcal{C}(\mathcal{P}_{1})$ for any best rank-one approximation.
@@
The following lemma will be applied in the next section to easily find a rank-two factorization. In the case of NMF this observation was used for hierarchical clustering in [29, Sec 2.3].

### v2

@@ 8 ND Rank One and Two
If ${\mathbf{T}}\in\mathcal{C}(\times_{j=1}^{k}\mathcal{P}_{j})$ is a rank-one tensor, then $\mathbf{T}$} has an ND rank of one.
@@
As ${\mathbf{T}}\in\mathcal{N}_{<\infty}$ we have that ${\mathbf{T}}=\sum_{i=1}^{r}{\mathbf{a}}^{(i)}\otimes{\mathbf{b}}^{(i)}$, ${\mathbf{a}}^{(i)}\in\mathcal{C}(\mathcal{P}_{1})$, ${\mathbf{b}}^{(i)}\in\mathcal{C}(\mathcal{P}_{2})$, and ${\mathbf{T}}{\mathbf{h}}=\sum_{i=1}^{r}({\mathbf{h}}^{\intercal}{\mathbf{b}}^{(i)}){\mathbf{a}}^{(i)}\in\mathcal{C}(\mathcal{P}_{1})$ for all ${\mathbf{h}}\in\mathcal{C}(\mathcal{P}_{2})^{*}$. Similarly, ${\mathbf{h}}^{\intercal}{\mathbf{T}}\in\mathcal{C}(\mathcal{P}_{2})$ for all ${\mathbf{h}}\in\mathcal{C}(\mathcal{P}_{1})^{*}$. As $\mathcal{C}(\mathcal{P}_{2})^{*}$ is full-dimensional [41] ${\mathbf{T}}(\mathcal{C}(\mathcal{P}_{2})^{*})\subseteq\mathcal{C}(\mathcal{P}_{1})$ is a two-dimensional cone. Let ${\mathbf{v}}^{(1)},{\mathbf{v}}^{(2)}$ be the extremal rays of the two-dimensional cone $\text{col}({\mathbf{T}})\cap\mathcal{C}(\mathcal{P}_{1})$. There exist ${\mathbf{w}}^{(1)},{\mathbf{w}}^{(2)}$ such that ${\mathbf{T}}={\mathbf{v}}^{(1)}\otimes{\mathbf{w}}^{(1)}+{\mathbf{v}}^{(2)}\otimes{\mathbf{w}}^{(2)}$ and it remains to show that ${\mathbf{w}}^{(i)}\in\mathcal{C}(\mathcal{P}_{2})$ for $i=1,2$. There must exist a face of $\mathcal{C}(\mathcal{P}_{1})$ that contains ${\mathbf{v}}^{(1)}$ but not ${\mathbf{v}}^{(2)}$, since otherwise $\lambda{\mathbf{v}}^{(1)}+(1-\lambda){\mathbf{v}}^{(2)}$ would be contained in $\mathcal{C}(\mathcal{P}_{1})$ for a small enough $\lambda>1$, contradicting the assumption that ${\mathbf{v}}^{(1)}$ was extremal in $\text{col}({\mathbf{T}})\cap\mathcal{C}(\mathcal{P}_{1})$. Let ${\mathbf{h}}\in\mathcal{C}(\mathcal{P}_{1})^{*}$ be a dual vector corresponding to the hyperplane that supports this face. Then ${\mathbf{h}}^{\intercal}{\mathbf{T}}=({\mathbf{h}}^{\intercal}{\mathbf{v}}^{(1)}){\mathbf{w}}^{(1)}+({\mathbf{h}}^{\intercal}{\mathbf{v}}^{(2)}){\mathbf{w}}^{(2)}=({\mathbf{h}}^{\intercal}{\mathbf{v}}^{(2)}){\mathbf{w}}^{(2)}\in\mathcal{C}(\mathcal{P}_{2})$, implying ${\mathbf{w}}^{(2)}\in\mathcal{C}(\mathcal{P}_{2})$ because ${\mathbf{h}}^{\intercal}{\mathbf{v}}^{(2)}>0$. The same argument shows that ${\mathbf{w}}^{(1)}\in\mathcal{C}(\mathcal{P}_{2})$. ∎
@@
A ramification of the above theorems is that defining equations for the semialgebraic sets [11, Def 2.1.4] of ND rank-one tensors and ND rank-two matrices are known. In the former case, a tensor $\mathbf{T}$} is in $\mathcal{N}_{\leq 1}$ if and only if it is monotonic and if it satisfies the determinantal equations for a rank-one tensor [34, Sec 3.4.1]. In the latter case, a matrix $\mathbf{T}$} is in $\mathcal{N}_{\leq 2}$ if and only if every $3\times 3$ minor of $\mathbf{T}$} vanishes and if $\mathbf{T}$} satisfies the inequalities for the $\mathcal{H}$-representation of $\mathcal{N}_{<\infty}$ described in Sections 5 and 6. Semialgebraic conditions under which a tensor of rank two also has a nonnegative rank of two are put forward in [3]. Whenever all order cones $\mathcal{C}(\mathcal{P}_{j})$ are simplicial such conditions can be easily translated, via Theorem 4, into conditions for a tensor to have an ND rank of two.
@@ 9 ND Border Rank Equals ND Rank
Due to the presence of noise, in many applications it is unlikely that a data tensor has an exact, low-rank, ND factorization. Instead a solution to an optimization problem of the form $\text{argmin}_{\bm{\theta}\in\mathcal{N}_{\leq r}}D({\mathbf{T}},\bm{\theta})$ is sought, where $D$ is a divergence that measures the discrepancy between the observed data tensor $\mathbf{T}$} and an approximating, low ND rank tensor $\bm{\theta}$. To ensure that a solution to this optimization problem exists it is important that the set $\mathcal{N}_{\leq r}$ be closed. The set of tensors with real tensor rank at most $r$ is not closed, which necessitates the introduction of the concept of border rank [34, Sec 2.4.5]. It is shown in this section, extending the proof for nonnegative tensor ranks provided in [38], that $\mathcal{N}_{\leq r}$ is closed and there is no need for the additional notion of border rank.
@@
Assume that $\bm{\theta}^{(n)}=\sum_{i=1}^{r}\lambda_{i}^{(n)}\otimes_{j=1}^{k}{\mathbf{v}}_{ij}^{(n)}$ is a sequence of tensors in $\mathcal{N}_{\leq r}$ with $\|{\mathbf{T}}-\bm{\theta}^{(n)}\|_{F}\rightarrow\inf_{\bm{\theta}\in\mathcal{N}_{\leq r}}\|{\mathbf{T}}-\bm{\theta}\|_{F}$. We will first show that $\bm{\theta}^{(n)}$ has a convergent subsequence that converges to a point in $\mathcal{N}_{\leq r}$. By the continuity of the objective function this point of convergence will be a solution to the given optimization problem. The ${\mathbf{v}}_{ij}^{(n)}\in\mathcal{C}(\mathcal{P}_{j})$ are assumed, without loss of generality, to be scaled so that $\|{\mathbf{v}}_{ij}^{(n)}\|_{2}=1$ and $\lambda_{i}^{(n)}\geq 0$. It is claimed that there exists an $M$ such that $\sup_{n}\lambda_{i}^{(n)}\leq M$ for all $i$. To show this, note that $\|{\mathbf{T}}-\bm{\theta}^{(n)}\|_{F}\geq\|\bm{\theta}^{(n)}\|_{F}-\|{\mathbf{T}}\|_{F}$ and
@@
where the last inequality follows because $\langle{\mathbf{v}}_{ij}^{(n)},{\mathbf{v}}_{i^{\prime}j}^{(n)}\rangle\geq 0$ for every $i,i^{\prime},j$ as $\mathcal{C}(\mathcal{P}_{j})\subseteq\mathbb{R}_{+}^{p_{j}}$. If $\lambda_{i}^{(n)}\rightarrow\infty$ for any $i$ then $\|{\mathbf{T}}-\bm{\theta}^{(n)}\|_{F}\rightarrow\infty$, a contradiction. It follows that there exists a convergent subsequence with $\lambda_{i}^{(n_{m})}\rightarrow\lambda_{i}$ and ${\mathbf{v}}_{ij}^{(n_{m})}\rightarrow{\mathbf{v}}_{ij}\in\mathcal{C}(\mathcal{P}_{j})$, with the latter inclusion following from $\mathcal{C}(\mathcal{P}_{j})$ being closed. Thus, $\bm{\theta}^{(n_{m})}\rightarrow\tilde{\bm{\theta}}=\sum_{i=1}^{r}\lambda_{i}\otimes_{j=1}^{k}{\mathbf{v}}_{ij}\in\mathcal{N}_{\leq r}$ is a convergent subsequence. The point of convergence $\tilde{\bm{\theta}}=\text{argmin}_{\bm{\theta}\in\mathcal{N}_{\leq r}}\|{\mathbf{T}}-\bm{\theta}\|_{F}$ is a solution to the optimization problem of interest. To show that $\mathcal{N}_{\leq r}$ is closed, first observe that if $\mathbf{T}$} is in the closure of $\mathcal{N}_{\leq r}$, then $\inf_{\bm{\theta}\in\mathcal{N}_{\leq r}}\|{\mathbf{T}}-\bm{\theta}\|_{F}=0$. However, the point $\tilde{\theta}\in\mathcal{N}_{\leq r}$, obtained from a convergent subsequence as described above, is a solution to the minimization problem, implying that $\|{\mathbf{T}}-\tilde{\bm{\theta}}\|_{F}=\inf_{\bm{\theta}\in\mathcal{N}_{\leq r}}\|{\mathbf{T}}-\bm{\theta}\|_{F}=0$. That is, ${\mathbf{T}}=\tilde{\bm{\theta}}\in\mathcal{N}_{\leq r}$. ∎
@@ 10 Finding Low ND Rank Approximations
Solving the optimization problem $\inf_{\bm{\theta}\in\mathcal{N}_{\leq r}}\|{\mathbf{T}}-\bm{\theta}\|_{F}$ introduced in the previous section is a one possibility for finding a low ND rank approximation to a data tensor $\mathbf{T}$}. However, when the entries of $\mathbf{T}$} consist of count data or positive data the Frobenius norm objective function may not be the most appropriate criteria to minimize. We examine a likelihood-based approach for estimating the mean of $\mathbf{T}$} in this section. It is assumed that the entries $T_{i_{1}\ldots i_{k}}$ are independently sampled from an exponential family of distributions with mean parameter [14, Ch 3] $E(T_{i_{1}\ldots i_{k}})=\theta_{i_{1}\ldots i_{k}}$ for every $(i_{1},\ldots,i_{k})\in\times_{j=1}^{k}[p_{j}]$, where $E(\cdot)$ is the expected value. The tensor $\bm{\theta}$ is constrained to have an ND rank of at most $r$ and is estimated by maximizing the observed likelihood
@@
The function $p({\mathbf{T}}|\bm{\theta})$ is the joint probability distribution or probability mass function of the tensor $\mathbf{T}$}. Different distributional assumptions about $T_{i_{1}\ldots i_{k}}$ lead to different choices of $p(T_{i_{1}\ldots i_{k}}|\theta_{i_{1}\ldots i_{k}})$ and hence different optimization problems. Of primary interest are the following standard distributions:
@@
    If $T_{i_{1}\ldots i_{k}}\in\mathbb{R}$ then we may assume that $T_{i_{1}\ldots i_{k}}\sim\mathcal{N}(\theta_{i_{1}\ldots i_{k}},\sigma^{2})$ are independent Gaussians. The corresponding optimization problem is
@@
    Note that the value of the variance $\sigma^{2}$ does not change the above optimization problem.
@@
These four optimization problems represent special cases of $\beta$-divergences, that are commonly applied to NMF problems [25], where $\beta$ is respectively equal to $2$, 11, and 0 under the distributional assumptions in $1.$ $3.$ and $4.$ above. Similar to Proposition 7.2 in [38], which shows that optimal nonnegative approximations exist for Bregman divergences, the next result shows that this also holds for nondecreasing factorizations, but does not require any additional constraints on $\bm{\theta}$.
@@
This follows for $1.$ by Theorem 12. For the remaining three problems it is seen that there exists an $M$ where any minimizer must be contained in the region $\mathcal{S}=\{\bm{\theta}:\theta_{i_{1}\ldots i_{k}}\leq M,\forall i_{1},\ldots,i_{k}\}$, as the objective functions in $3.$ and $4.$ diverge to $\infty$ whenever $\theta_{i_{1}\ldots i_{k}}\rightarrow\infty$. In case $2.$ $\bm{\theta}$ is trivially contained in such an $\mathcal{S}$ because the probability simplex is compact. As $\mathcal{S}\cap\mathcal{N}_{\leq r}$ is compact by Theorem 12 and the objective functions are continuous on this region, existence of the solutions follows. Here the objective functions are extended continuously to take values in $\mathbb{R}\cup\{\infty\}$ with $T_{i_{1}\ldots i_{k}}\log(0)\coloneq\infty$ and $T_{i_{1}\ldots i_{k}}/0\coloneq\infty$ when $T_{i_{1}\ldots i_{k}}\neq 0$. ∎
@@
The Perron-Frobenius theorem [8, Thm 3.2] that states that nonnegative, square matrices have nonnegative leading eigenvectors, has been extended to nonnegative, cubical tensors in [37, Thm 1]. The next result is reminiscent of the Perron-Frobenius theorem, but is instead concerned with the singular vectors and values of a matrix rather than the eigenvectors and eigenvalues.
@@
If every fibre ${\mathbf{T}}_{i_{1}\ldots i_{j-1}\bullet i_{j+1}\ldots i_{k}}$ is in $\mathcal{C}(\mathcal{P}_{j})$ then the best rank-one approximation to a non-zero $\mathbf{T}$} with respect to any of the log-likelihoods mentioned in this section is equal to the best ND rank-one approximation:
$\underset{\bm{\theta}:\mathrm{rank}(\bm{\theta})=1}{\mathrm{argmax}}\;p({\mathbf{T}}|\bm{\theta})=\underset{\bm{\theta}\in\mathcal{N}_{1}}{\mathrm{argmax}}\;\;p({\mathbf{T}}|\bm{\theta}).$
@@
Let $\bm{\theta}=\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}$ be a rank-one tensor, where we seek vectors ${\mathbf{v}}^{(j)}\in\mathbb{R}^{p_{j}}$ that maximize the likelihood of interest. Under the Gaussian likelihood, upon taking a gradient with respect to ${\mathbf{v}}^{(j)}$, the first-order conditions, without any monotonicity constraints, for $\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}$ to be optimal are
@@
We may assume that any optimal rank-one approximation has ${\mathbf{v}}^{(j)}\geq{\mathbf{0}}$ as the tensor $\max(\otimes_{j=1}^{k}{\mathbf{v}}^{(j)},{\mathbf{0}})=\otimes_{j=1}^{k}\max({\mathbf{v}}^{(j)},{\mathbf{0}})$, where the maximum is taken elementwise, is closer to $\mathbf{T}$} than $\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}$ is when $\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}$ has a negative entry. The optimal ${\mathbf{v}}^{(j)}$ is non-zero when $\mathbf{T}$} is non-zero since if $T_{i_{1}\ldots i_{k}}\neq 0$ then $\bm{\theta}=T_{i_{1}\ldots i_{k}}\otimes_{j=1}^{k}{\mathbf{e}}_{i_{j}}$ has a larger likelihood value than $\bm{\theta}={\mathbf{0}}$. Equation (10) shows that ${\mathbf{v}}^{(j)}$ must be a conic combination of the monotone fibres ${\mathbf{T}}_{i_{1}\ldots i_{j-1}\bullet i_{j+1}\ldots i_{k}}$, and so it must itself be in $\mathcal{C}(\mathcal{P}_{j})$.
When $k=2$ and ${\mathbf{v}}^{(1)},{\mathbf{v}}^{(2)}$ are any choice of the first left and right singular vectors, by the Eckhart-Young theorem, ${\mathbf{v}}^{(1)}\otimes{\mathbf{v}}^{(2)}$ is a best rank-one approximation of $\mathbf{T}$}. Therefore, by the above derivation, ${\mathbf{v}}^{(j)}\in\mathcal{C}(\mathcal{P}_{j})$, up to a possible sign change. If the largest singular value was repeated then we have pairs ${\mathbf{v}}^{(j)},{\mathbf{w}}^{(j)}$ of left ($j=1$) and right ($j=2$) singular vectors. As $\mathcal{C}(\mathcal{P}_{1})$ is proper, the subspace $\text{span}({\mathbf{v}}^{(1)},{\mathbf{w}}^{(1)})$ contains a unit norm vector ${\mathbf{z}}^{(1)}\notin\pm\mathcal{C}(\mathcal{P}_{1})$. Defining the $p_{j}\times 2$ matrix ${\mathbf{R}}^{(j)}=[{\mathbf{v}}^{(j)}|{\mathbf{w}}^{(j)}]$, consider the rotation matrix $\mathbf{V}$} constructed so that the first column of ${\mathbf{R}}^{(1)}{\mathbf{V}}$ is equal to ${\mathbf{z}}^{(1)}$. Set ${\mathbf{z}}^{(2)}$ to be the first column of ${\mathbf{R}}^{(2)}{\mathbf{V}}$. We have that ${\mathbf{R}}^{(1)}({\mathbf{R}}^{(2)})^{\intercal}={\mathbf{R}}^{(1)}{\mathbf{V}}{\mathbf{V}}^{\intercal}({\mathbf{R}}^{(2)})^{\intercal}$ is a best rank-two approximation of $\mathbf{T}$} and thus ${\mathbf{z}}^{(1)}\otimes{\mathbf{z}}^{(2)}$ is a best rank-one approximation of $\mathbf{T}$}. This contradicts the above fact that ${\mathbf{z}}^{(1)}\in\pm\mathcal{C}(\mathcal{P}_{1})$ for any best rank-one approximation.
@@
The following lemma will be applied in the next section to easily find a rank-two factorization. In the case of NMF this observation was used for hierarchical clustering in [30, Sec 2.3].